What is the mass of a bus moving at a velocity of 110 kph if it has to fully stop before an intersection if a 7000-N of force is applied on it in 15 seconds? *


1
Expert's answer
2022-04-06T13:57:58-0400

Given:

"\Delta v=110\:\rm kph=30.6\: m/s"

"F=7000\:\rm N"

"\Delta t=15\:\rm s"


The Newton's second law says

"F=\frac{m\Delta v}{\Delta t}"

Hence, the mass of a bus

"m=\frac{F\Delta t}{\Delta v}=\frac{7000*15}{30.6}=3436\:\rm kg"